# Fractal counterparts of density combinatorics theorems in continued fractions

### v1

@@ 4. Proofs of the main results
Let $S$ be an infinite subset of $\mathbb{N}$. By Lemma 2.7, if $x\in E$ then $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)>0$ implies $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=1$ which in turn implies $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})=1$. Applying Lemma 2.8 with $c=1$ we conclude that $\left\{x\in J\colon\ \overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)>0\right\}$ is of Hausdorff dimension zero. So, if $\overline{d}(S)>0$ then $\left\{x\in J\colon\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)\right\}$ is of Hausdorff dimension zero. Since $\dim_{\rm H}J=1/2>0$ by Lemma 3.4, the desired inequality in Theorem 1.13(a) holds.
@@
Suppose $S$ has polynomial density with exponent $\alpha<2$. Lemma 2.7 gives $\tau(S)=1/\alpha$. Let $A\subset S$ satisfy $\overline{d}(A|S)>0$. By Lemma 2.7, if $x\in E$ then $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)>0$ implies $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A)=1/\alpha$ which in turn implies $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})\geq 1/\alpha$. If moreover $\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S$ then $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})\leq\tau(S)=1/\alpha$. Further, applying Lemma 2.8 with $c=1/\alpha$ we obtain
$\dim_{\rm H}\left\{x\in J\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S,\ \overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)>0\right\}\leq\frac{\alpha-1}{2\alpha}.$
@@
Combining this inequality, Corollary 3.4 and then using $\alpha<2$ yields the desired inequality in Theorem 1.13(b).∎
@@ Appendix A Bergelson-Leibman’s theorem in $\mathbb{N}$
Let $\sigma$ denote the left shift acting on $\{0,1\}^{\mathbb{Z}}$ and let $\Sigma=\overline{\{\sigma^{i}(x^{*})\colon i\in\mathbb{Z}\}}$. Then $\Sigma$ is a compact metrizable space and $\sigma(\Sigma)=\Sigma$. Let $A=\{(y_{i})_{i=-\infty}^{\infty}\in\Sigma\colon y_{0}=1\}$. Let $(I_{k})_{k=1}^{\infty}$ be a sequence of compact intervals in $[1,\infty)$ such that #$\#(\mathbb{N}\cap I_{k})\to\infty$ as $k\to\infty$ and
@@
For each $k\geq 1$ define a Borel probability measure $\mu_{k}$ on $\Sigma$ by
$\mu_{k}=\frac{1}{\#(\mathbb{N}\cap I_{k})}\sum_{i\in\mathbb{N}\cap I_{k}}\delta_{\sigma^{i}(x^{*})},$
where $\delta_{\sigma^{i}(x^{*})}$ denotes the unit point mass at $\sigma^{i}(x^{*})$. Note that $\lim_{k\to\infty}\mu_{k}(A)=d_{\rm B}(S)$. Pick a limit point of the sequence $(\mu_{k})_{k=1}^{\infty}$ in the weak* topology and denote it by $\mu$. Then $\mu$ is $\sigma|_{\Sigma}$-invariant. Since $A$ is a clopen set, we have $\mu(A)=\lim_{k\to\infty}\mu_{k}(A)$ and so $\mu(A)=d_{\rm B}(S)>0$.
For each $m\in\mathbb{Z}$ let $S-m=\{i\in\mathbb{N}\colon i+m\in S\}.$ Let $k\geq 1$ and $m_{1},\ldots,m_{\ell}\in\mathbb{Z}$. For each $i\in\mathbb{N}$, note that $i\in I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell})$ if and only if $\sigma^{i}(x^{*})\in A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A)$. This correspondence yields
#$\begin{split}\frac{\#(I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))}{\#(\mathbb{N}\cap I_{k})}=\mu_{k}(A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A)).\end{split}$
@@
$\begin{split}d_{\rm B}(S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))&\geq\lim_{k\to\infty}\frac{\#(I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))}{\#(\mathbb{N}\cap I_{k})}\\ &=\lim_{k\to\infty}\mu_{k}(A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A))\\ &=\mu(A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A)).\end{split}$
The first inequality follows from the definition of upper Banach density. The second equality holds because $A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A)$ is a clopen set. Let $h_{1},\ldots,h_{\ell}\in\mathbb{Z}_{0}[X]$. Replacing $m_{j}$ by $h_{j}(m)$, $m\in\mathbb{N}$ for $j=1,\ldots,\ell$ we have
$d_{\rm B}(S\cap(S-h_{1}(m))\cap\cdots\cap(S-h_{\ell}(m)))\geq\mu(A\cap\sigma^{-h_{1}(m)}(A)\cap\cdots\cap\sigma^{-h_{\ell}(m)}(A)).$
Applying Theorem A.1 to the invertible $\mu$-preserving system $(\Sigma,\sigma|_{\Sigma})$ we obtain $d_{\rm B}(S\cap(S-h_{1}(m))\cap\cdots\cap(S-h_{\ell}(m)))>0$ for infinitely many $m\in\mathbb{N}$. This yields the desired statement.∎

### v2

@@ 4. Proofs of the main results
Let $S$ be an infinite subset of $\mathbb{N}$. By Lemma 2.7, if $x\in E$ then $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)>0$ implies
$\tau(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=1=\tau(\{a_{n}(x)\colon n\in\mathbb{N}\}).$
Applying Lemma 2.9 with $c=1$, we conclude that
$\dim_{\rm H}\left\{x\in J\colon\ \overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)>0\right\}=0.$
Since $\dim_{\rm H}J=1/2>0$ by [35] or Proposition 3.4, the desired inequality in Theorem 1.13(a) holds.
@@
Suppose $S$ has polynomial density with exponent $\alpha<2$. Lemma 2.7 gives $\tau(S)=1/\alpha$. Let $A\subset S$ satisfy $\overline{d}(A|S)>0$. By Lemma 2.7, if $x\in E$ then $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)>0$ implies
$\tau(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A)=\frac{1}{\alpha}\leq\tau(\{a_{n}(x)\colon n\in\mathbb{N}\}).$
If moreover $\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S$, then we have
$\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})\leq\tau(S)=\frac{1}{\alpha}.$
Further, applying Lemma 2.9 with $c=1/\alpha$ we obtain
[…]
dimH{ :xJ
@@
Combining this inequality, the equality in Proposition 3.4 and then using $\alpha<2$ yields the desired inequality in Theorem 1.13(b).∎
5. Some Generalizations
In this last section, we generalize Theorems 1.1 and 1.4 to other expansions of real numbers. Proofs are essentially the same as that of Theorems 1.1 and 1.4, and hence will be omitted.
5.1. Iterated Function Systems
A countable collection of self-maps is called an Iterated Function System (IFS) [8]. Let $\Phi=\{\varphi_{k}\colon k\in\mathbb{N}\}$ be a collection of $C^{1}$ maps from $[0,1]$ to itself and let $d>1$. We say $\Phi$ is a $d$-decaying IFS [19] if the following conditions hold:
  • (i)
    there exists $m\in\mathbb{N}$ and $\gamma\in(0,1)$ such that for every $(a_{1},\ldots,a_{m})\in\mathbb{N}^{m}$ and every $x\in[0,1]$ we have $0<|(\varphi_{a_{1}}\circ\cdots\circ\varphi_{a_{m}})^{\prime}(x)|\leq\gamma;$
  • (ii)
    for all $i,j\in\mathbb{N}$ with $i\neq j$, $\varphi_{i}((0,1))\cap\varphi_{j}((0,1))=\emptyset;$
  • (iii)
    for any $\varepsilon>0,$ there exist positive constants $c_{1}(\varepsilon)$, $c_{2}(\varepsilon)$ such that for every $n\in\mathbb{N}$ and every $x\in[0,1]$ we have
    $\frac{c_{1}(\varepsilon)}{k^{d(1+\varepsilon)}}\leq|\varphi_{k}^{\prime}(x)|\leq\frac{c_{2}(\varepsilon)}{k^{d(1-\varepsilon)}};$
  • (iv)
    $\overline{\bigcup_{k=1}^{\infty}\varphi_{k}([0,1])}=[0,1]$, and if $i<j$ then $\varphi_{i}(x)>\varphi_{j}(x)$ for all $x\in[0,1].$
If $\Phi=\{\varphi_{k}\colon k\in\mathbb{N}\}$ is a $d$-decaying IFS, then by (i) one can define a natural projection $\Pi\colon\mathbb{N}^{\mathbb{N}}\rightarrow[0,1]$ by
$\Pi((a_{n})_{n=1}^{\infty})=\lim_{n\to\infty}\varphi_{a_{1}}\circ\cdots\circ\varphi_{a_{n}}(1).$
By (ii), $\Pi$ is one-to-one except on a countable number of points. If $x\in[0,1]$ and $\Pi^{-1}(x)$ is a singleton, then there exists a unique sequence $(a_{n}(x))_{n=1}^{\infty}$ of positive integers such that
$x=\Pi((a_{n}(x))_{n=1}^{\infty}).$
This equation may be viewed as an expansion of $x$ with the digit sequence $(a_{n}(x))_{n=1}^{\infty}$. As an analogue of the set $E$ for the regular continued fraction, define $E(\Phi)$ to be the set of $x\in[0,1]$ such that $\Pi^{-1}(x)$ is a singleton and $a_{m}(x)\neq a_{n}(x)$ holds for all $m$, $n\in\mathbb{N}$ with $m\neq n$. Theorems 1.1 and 1.4 can be generalized to expansions generated by arbitrary $d$-decaying IFSs as follows.
Theorem 5.1.
Let $d>1$ and let $\Phi$ be a $d$-decaying IFS. Let $S\subset\mathbb{N}$.
  • (a)
    If $d_{\rm B}(S)>0$, then there exists $F_{S}(\Phi)\subset E(\Phi)$ such that
    $\dim_{\rm H}F_{S}(\Phi)=\dim_{\rm H}E(\Phi)\ \text{ and }\ d_{\rm B}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}(\Phi)}\{a_{n}(x)\}\cap S\right)=d_{\rm B}(S).$
    In particular,
    $\dim_{\rm H}\left\{x\in E(\Phi)\colon d_{\rm B}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=d_{\rm B}(S)\right\}=\dim_{\rm H}E(\Phi).$
  • (b)
    If $S$ has polynomial density with exponent $\alpha\geq 1$, then for any $A\subset S$ with $\overline{d}(A|S)>0$ there exists a subset $E_{S,A}(\Phi)$ of $\{x\in E(\Phi)\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}$ such that
    $\dim_{\rm H}E_{S,A}(\Phi)=\dim_{\rm H}\{x\in E(\Phi)\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}=\frac{1}{d\alpha}$
    and
    $\overline{d}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}(\Phi)}\{a_{n}(x)\}\cap A|S\right)=\overline{d}(A|S).$
    In particular,
[…]
    dimH{ :xE(Φ)
The regular continued fraction is generated by the $2$-decaying IFS $\{\varphi_{k}\colon k\in\mathbb{N}\}$ defined by $\varphi_{k}(x)=1/(x+k)$ with $m=2$. Hence, Theorem 5.1 is indeed a generalization of Theorems 1.1 and 1.4. Then, Theorems 1.3, 1.5, 1.6, 1.9, 1.11 and 1.12 for the regular continued fraction are generalized to $d$-decaying IFSs accordingly. Precise statements are omitted.
Regarding a proof of Theorem 5.1, most of the arguments in the proofs of Theorems 1.1 and 1.4 remain intact. Lemma 2.1 is the only place where the specific formulas for the regular continued fraction were used. For each $n\in\mathbb{N}$, the intervals $\varphi_{a_{1}}\circ\cdots\circ\varphi_{a_{n}}([0,1])$, $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ play the role of the $n$-th fundamental intervals. As a substitute for Lemma 2.1, one can use (iii) to evaluate the diameters of these intervals.
5.2. Semi-regular continued fractions
Let $\sigma=\{\sigma_{n}\}_{n=1}^{\infty}\in\{-1,1\}^{\mathbb{N}}$. Any number $x$ in the interval $(0,1)$ has the unique, finite or infinite continued fraction expansion of the form
$x=\frac{\displaystyle{\hfill{1}\hfill\;\vrule}}{\displaystyle{\vrule\;\hfill{a_{\sigma,1}(x)}\hfill}}+\frac{\displaystyle{\hfill{\sigma_{1}}\hfill\;\vrule}}{\displaystyle{\vrule\;\hfill{a_{\sigma,2}(x)}\hfill}}+\frac{\displaystyle{\hfill{\sigma_{2}}\hfill\;\vrule}}{\displaystyle{\vrule\;\hfill{a_{\sigma,3}(x)}\hfill}}+\cdots,$ (5.1)
where $a_{\sigma,n}(x)$ are positive integers such that $\sigma_{n}+a_{\sigma,n}(x)\geq 1$ for every $n\geq 1$ [20, 30]. This kind of continued fractions are called semi-regular [17, 22]. If $x$ is irrational then the continued fraction in (5.1) is infinite, which means that the finite truncation
$\cfrac{1}{a_{\sigma,1}(x)+\cfrac{\sigma_{1}}{a_{\sigma,2}(x)+\cdots+\cfrac{\ddots}{\displaystyle{\frac{\sigma_{n-1}}{a_{\sigma,n}(x)}}}}}$
converges to $x$ as $n\to\infty$. If $\sigma_{n}=1$ for all $n\in\mathbb{N}$ (resp. $\sigma_{n}=-1$ for all $n\in\mathbb{N}$) then we get the regular (resp. backward) continued fraction expansion. For other examples and motivations for semi-regular continued fractions, see [17, 22] and the references therein.
As an analogue of the set $E$ for the regular continued fraction, for each $\sigma\in\{-1,1\}^{\mathbb{N}}$ let $E_{\sigma}$ denote the set of $x\in(0,1)\setminus\mathbb{Q}$ such that $a_{\sigma,m}(x)\neq a_{\sigma,n}(x)$ holds for all $m$ $n\in\mathbb{N}$ with $m\neq n$. Theorems 1.1 and 1.4 can be generalized to semi-regular continued fractions as follows.
Theorem 5.2.
Let $S\subset\mathbb{N}$.
  • (a)
    If $d_{\rm B}(S)>0$, then for any $\sigma\in\{-1,1\}^{\mathbb{N}}$ there exists $F_{S,\sigma}\subset E_{\sigma}$ such that
    $\dim_{\rm H}F_{S,\sigma}=\dim_{\rm H}E_{\sigma}\ \text{ and }\ d_{\rm B}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S,\sigma}}\{a_{n,\sigma}(x)\}\cap S\right)=d_{\rm B}(S).$
    In particular,
    $\dim_{\rm H}\left\{x\in E_{\sigma}\colon d_{\rm B}(\{a_{n,\sigma}(x)\colon n\in\mathbb{N}\}\cap S)=d_{\rm B}(S)\right\}=\dim_{\rm H}E_{\sigma}.$
  • (b)
    If $S$ has polynomial density with exponent $\alpha\geq 1$, then for any $A\subset S$ with $\overline{d}(A|S)>0$ and any $\sigma\in\{-1,1\}^{\mathbb{N}}$, there exists a subset $E_{S,A,\sigma}$ of $\{x\in E_{\sigma}\colon\{a_{\sigma,n}(x)\colon n\in\mathbb{N}\}\subset S\}$ such that
    $\dim_{\rm H}E_{S,A,\sigma}=\dim_{\rm H}\{x\in E_{\sigma}\colon\{a_{\sigma,n}(x)\colon n\in\mathbb{N}\}\subset S\}=\frac{1}{2\alpha}$
    and
    $\overline{d}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A,\sigma}}\{a_{\sigma,n}(x)\}\cap A|S\right)=\overline{d}(A|S).$
    In particular,
[…]
    dimH{ :xEσ
By virtue of Theorem 5.2, Theorems 1.3, 1.5, 1.6, 1.9, 1.11 and 1.12 are generalized to semi-regular continued fractions accordingly. Precise statements are omitted.
Regarding a proof of Theorem 5.2, most of the arguments in the proofs of Theorems 1.1 and 1.4 remain intact apart from Lemma 2.1. For each $\sigma\in\{-1,1\}^{\mathbb{N}}$, $n\in\mathbb{N}$ and $(a_{1},\ldots,a_{n})\in\mathbb{N}$ such that $\sigma_{i}+a_{i}\geq 1$ for $i=1,\ldots,n$, we define
$I_{\sigma}(a_{1},\ldots,a_{n})=\{x\in(0,1)\colon a_{\sigma,i}(x)=a_{i}\ \text{ for }i=1,\ldots,n\}.$
These sets are non-degenerate intervals, and play the role of the $n$-th fundamental intervals. The following can be used as a substitute for Lemma 2.1.
Lemma 5.3.
There exists a constant $C\geq 1$ such that if $\sigma\in\{-1,1\}^{\mathbb{N}}$, $n\in\mathbb{N}$, $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ are such that $\sigma_{i}+a_{i}\geq 1$ and $a_{i}\geq 3$ for $i=1,\ldots,n$ then
$C^{-1}\prod_{i=1}^{n}\frac{1}{(a_{i}+1)^{2}}\leq|I_{\sigma}(a_{1},\ldots,a_{n})|\leq C\prod_{i=1}^{n}\frac{1}{(a_{i}-1)^{2}}.$
Similarly to the case of the regular continued fraction, the endpoints of the intervals $I_{\sigma}(a_{1},\ldots,a_{n})$ can be written with the partial numerators and denominators of the semi-regular continued fraction expansion (see [22]). Hence one can prove Lemma 5.3 similarly to the proof of Lemma 2.1. Here we give a conceptually clearer, dynamical alternative proof.
Proof of Lemma 5.3.
Define two maps $T_{1}\colon x\in(0,1]\mapsto 1/x-\lfloor 1/x\rfloor$ and $T_{-1}\colon x\in(0,1]\mapsto\lfloor 1/x\rfloor-1/x+1$. For each $k\in\mathbb{N}$, $T_{1}$ and $T_{-1}$ send the interval $(\frac{1}{k+1},\frac{1}{k})$ diffeomorphically onto $(0,1)$. They satisfy Rényi’s condition
[…]
Write $T_{\sigma}^{n}=T_{\sigma_{n}}\circ\cdots\circ T_{\sigma_{1}}$, and for convenience set $T_{\sigma}^{0}$ to be the identity map on $(0,1)$. Let $\theta$ denote the left shift acting on $\{-1,1\}^{\mathbb{N}}$. We have $T_{\sigma}^{i}(I_{\sigma}(a_{1},\ldots,a_{n}))=I_{\theta^{i}\sigma}(a_{i+1},\ldots,a_{n})$ for $i=1,\ldots,n-1$, and $T_{\sigma}^{n}$ sends $I_{\sigma}(a_{1},\ldots,a_{n})$ diffeomorphically onto $(0,1)$. For all $x\in I_{\sigma}(a_{1},\ldots,a_{n})$ and $i=1,\ldots,n$ we have
$(a_{i}-1)^{2}\leq|(T_{\sigma_{i}})^{\prime}T_{\sigma}^{i-1}x|\leq(a_{i}+1)^{2}.$ (5.2)
Since $a_{1},\ldots,a_{n}\geq 3$, combining the first inequality in (5.2) and Rényi’s condition we obtain a constant $C\geq 1$ independent of $\sigma$, $n$ such that
$\sup_{x,y\in I_{\sigma}(a_{1},\ldots,a_{n})}\frac{|(T_{\sigma}^{n})^{\prime}y|}{|(T_{\sigma}^{n})^{\prime}x|}\leq C.$
By the mean value theorem, there exists $x\in I_{\sigma}(a_{1},\ldots,a_{n})$ such that $C^{-1}|(T_{\sigma}^{n})^{\prime}x|^{-1}\leq|I_{\sigma}(a_{1},\ldots,a_{n})|\leq C|(T_{\sigma}^{n})^{\prime}x|^{-1}$. The chain rule gives $|(T^{n}_{\sigma})^{\prime}x|=\prod_{i=1}^{n}|(T_{\sigma_{i}})^{\prime}T_{\sigma}^{i-1}x|$. Using (5.2) we obtain the desired double inequalities. ∎
@@ Appendix A Bergelson-Leibman’s theorem in $\mathbb{N}$
Let $\theta$ denote the left shift acting on $\{0,1\}^{\mathbb{Z}}$ and let $\Sigma=\overline{\{\theta^{i}(x^{*})\colon i\in\mathbb{Z}\}}$. Then $\Sigma$ is a compact metrizable space and $\theta(\Sigma)=\Sigma$. Let $A=\{(y_{i})_{i=-\infty}^{\infty}\in\Sigma\colon y_{0}=1\}$. Let $(I_{k})_{k=1}^{\infty}$ be a sequence of compact intervals in $1,\infty)$ such that $\lim_{k\to\infty}\#(\mathbb{N}\cap I_{k})=\infty$ and
@@
For each $k\geq 1$, define a Borel probability measure $\mu_{k}$ on $\Sigma$ by
$\mu_{k}=\frac{1}{\#(\mathbb{N}\cap I_{k})}\sum_{i\in\mathbb{N}\cap I_{k}}\delta_{\theta^{i}(x^{*})},$
where $\delta_{\theta^{i}(x^{*})}$ denotes the unit point mass at $\theta^{i}(x^{*})$. Note that $\lim_{k\to\infty}\mu_{k}(A)=d_{\rm B}(S)$. Pick a limit point of the sequence $(\mu_{k})_{k=1}^{\infty}$ in the weak* topology and denote it by $\mu$. Then $\mu$ is $\theta|_{\Sigma}$-invariant. Since $A$ is a clopen set, we have $\mu(A)=\lim_{k\to\infty}\mu_{k}(A)$ and so $\mu(A)=d_{\rm B}(S)>0$.
For each $m\in\mathbb{Z}$ define
$S-m=\{i\in\mathbb{N}\colon i+m\in S\}.$
Let $k\geq 1$ and $m_{1},\ldots,m_{\ell}\in\mathbb{Z}$. Note that $i\in I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell})$ if and only if $\theta^{i}(x^{*})\in A\cap\theta^{-m_{1}}(A)\cap\cdots\cap\theta^{-m_{\ell}}(A)$. This correspondence yields
$\begin{split}\frac{\#(I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))}{\#(\mathbb{N}\cap I_{k})}=\mu_{k}(A\cap\theta^{-m_{1}}(A)\cap\cdots\cap\theta^{-m_{\ell}}(A)).\end{split}$
@@
$\begin{split}d_{\rm B}(S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))&\geq\lim_{k\to\infty}\frac{\#(I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))}{\#(\mathbb{N}\cap I_{k})}\\ &=\lim_{k\to\infty}\mu_{k}(A\cap\theta^{-m_{1}}(A)\cap\cdots\cap\theta^{-m_{\ell}}(A))\\ &=\mu(A\cap\theta^{-m_{1}}(A)\cap\cdots\cap\theta^{-m_{\ell}}(A)).\end{split}$
The first inequality follows from the definition of upper Banach density. The second equality holds because $A\cap\theta^{-m_{1}}(A)\cap\cdots\cap\theta^{-m_{\ell}}(A)$ is a clopen set. Let $h_{1},\ldots,h_{\ell}\in\mathbb{Z}_{0}[X]$. Replacing $m_{j}$ by $h_{j}(m)$, $m\in\mathbb{N}$ for $j=1,\ldots,\ell$ we have
$d_{\rm B}(S\cap(S-h_{1}(m))\cap\cdots\cap(S-h_{\ell}(m)))\geq\mu(A\cap\theta^{-h_{1}(m)}(A)\cap\cdots\cap\theta^{-h_{\ell}(m)}(A)).$
Applying Theorem A.1 to the invertible $\mu$-preserving system $(\Sigma,\theta|_{\Sigma})$, we obtain
$d_{\rm B}(S\cap(S-h_{1}(m))\cap\cdots\cap(S-h_{\ell}(m)))>0$
for infinitely many $m\in\mathbb{N}$. This yields the desired statement.∎